OpenStudy (anonymous):

the integral of csc^4 (x) cot^6 (x) dx

OpenStudy (anonymous):

You know that the derivative of cotangent is negative cosecant squared, so I'd start off with breaking up the powers like so: \[\csc^4x\cot^6x=\csc^2x\csc^2x\cot^6x\] Now apply a trig identity to one of the csc terms: \[\cdots=\csc^2x\left(1+\cot^2x\right)\cot^6x\] Then distribute: \[\cdots=\csc^2x\left(\cot^6x+\cot^8x\right)\] So to integrate \(\displaystyle\int \csc^2x\left(\cot^6x+\cot^8x\right)~dx\), substitute \(u=\cot x\).

OpenStudy (anonymous):

I may have gotten it \[-\frac{ \cot ^{7} }{ 7 } - \frac{ \cot ^{9}}{ 9 }\] +constant

OpenStudy (anonymous):

Yes, that's correct

OpenStudy (anonymous):

Oh awesome! thank you for the help!

OpenStudy (anonymous):

yw